OpenStudy (anonymous):

integrate 7x cos^4(x^2) dx

OpenStudy (anonymous):

Try \[ u=x^2\quad du=2x\;dx \]

OpenStudy (anonymous):

Also \[ \cos^2(u) = \frac{1+\cos(2u)}{2} \]

OpenStudy (anonymous):

The only question is, how badly do you want it?

OpenStudy (anonymous):

The 7x is confusing me

OpenStudy (anonymous):

My instinct says to solve it like an integration by parts

OpenStudy (anonymous):

Okay, here is some basic algebra: \(7x=3.5\cdot 2x\)

OpenStudy (anonymous):

\[ 7x = \frac 7 2 \cdot 2x \]

OpenStudy (anonymous):

\[ \int 7x \cos^4(x^2) dx = \int \frac 7 2 \cos^4(x^2)\;2xdx \]

OpenStudy (anonymous):

It's not like there is any obvious pattern here or anything.\[ u = x^2\quad du=2xdx \]

OpenStudy (anonymous):

ah, okay

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 7 }{ 2 }\cos ^{4}(u) du\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio Also \[ \cos^2(u) = \frac{1+\cos(2u)}{2} \] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

so you plug that in for cos and square the whole formula?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

and bring out the 7/2 constant

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

\[\frac{ 7 }{ 2 }\int\limits_{}^{}(\frac{ 1+\cos(2u) }{ 1 })^{2}u du\]

OpenStudy (anonymous):

Not \(udu\), just \(du\)

OpenStudy (anonymous):

2 in the denominator

OpenStudy (anonymous):

Now use foil.

OpenStudy (anonymous):

what happened to the x^2 in cos^4x^2?

OpenStudy (anonymous):

\(u=x^2\)

OpenStudy (anonymous):

I'm very confused for some reason. Is the half angle formula squared because of the x^2 or the cos^4?

OpenStudy (anonymous):

I was thinking that it was squared because of the 4th root on cos. That's why I put u du on the end and not just du

OpenStudy (anonymous):

It's because the \(\cos\) is squared.

OpenStudy (anonymous):

oh, I think it just clicked

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[ \cos^4(u) = [\cos^2(u)]^2 = \left[\frac{1+\cos(2u)}{2}\right]^2 \]

OpenStudy (anonymous):

so now we're multiplying it out

OpenStudy (anonymous):

\[\frac{ 7 }{ 2 }\int\limits_{}^{}\frac{ 1+2\cos(2u)+\cos ^{2(2u)} }{ 4} du\]

OpenStudy (anonymous):

not an exponent on the "2u" obviously

OpenStudy (anonymous):

Pull out the 4 and use the half angle again on the squared cosine. Then it will be simple.

OpenStudy (anonymous):

okay, I got this

OpenStudy (anonymous):

I doubt I did this right\[\frac{ 14 }{ 16 }\int\limits_{}^{}2+2\cos(2u) du\]

OpenStudy (anonymous):

Show your work.

OpenStudy (anonymous):

\[\frac{ 7 }{ 8 }\int\limits_{}^{} 1+2\cos(2u)+(\frac{ 1+\cos(2u) }{ 2 }) du\]

OpenStudy (anonymous):

should be \(4u\)

OpenStudy (anonymous):

on the first cos?

OpenStudy (anonymous):

Okay, which one do you think I was talking about?

OpenStudy (anonymous):

the first one. I didn't have to do anything with the second besides plug in the formula

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

I dont see why either one would be 4u

OpenStudy (anonymous):

and I wouldn't have asked in the first place if I knew

OpenStudy (anonymous):

\[ \cos^2(2u) = \frac{1+\cos(4u)}{2} \]